OpenStudy (anonymous):

Write the equation in slope-intercept form. Then tell whether its graph has a positive or negative slope. a. y=-3+8; negative slope c. y=3x-8; positive slope b. y=3x+8; positive slope d. y=3x+8: negative slope

OpenStudy (anonymous):

http://student.education2020.com/Calculators/graphing-calculator.html here is a graph of that helps??

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What equation?

OpenStudy (anonymous):

-3x+y=8

OpenStudy (anonymous):

Slope intercept form simply means write it like this y=mx+b

OpenStudy (anonymous):

so it is either b or d?

OpenStudy (anonymous):

You don't have to guess. You just have to move -3x to the other side.

OpenStudy (anonymous):

y=-3x+b

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do you mean y=-3+8 @kandsbbf

OpenStudy (anonymous):

yes, sorry about that

OpenStudy (anonymous):

it is okay thx

OpenStudy (anonymous):

what is the answer b or d ??? idk

OpenStudy (anonymous):

not c or a

OpenStudy (anonymous):

You are confused because you and kandsbbf didn't work the equation correctly.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

how didn't i it's asking for the equation in y=mx+b

OpenStudy (anonymous):

Let's try again, write -3x + y = 8 in ...

OpenStudy (anonymous):

ok